You take out a 30 year £200,000 mortgage loan with monthly payments and a 0.5% monthly interest rate. Fourteen years later you decide to sell your house and pay off the mortgage. What is the balance on the loan?
1
Expert's answer
2019-12-17T09:50:16-0500

The monthly payment is:

"M = 200,000×\frac{0.005}{1 - (1 + 0.005)^{-360}} = £1199.1."

The sum to be repaid in 14 years is:

1199.1×14×12 = £230,227.2.

If you decide to sell your house and pay off the mortgage, then the balance of the loan will be: 230,227.2 - 200,000 = 30,227.2 to be repaid.